OpenStudy (anonymous):

A fan club that comprises of boys and girls have a total of 160 members. If 32 more girls join the club, the percentage of girls will increase by 5%. Find the number of boys in the club.

mathslover (mathslover):

ok ! total members = 160 let the no. of girls = x and no. of boys = y ok ?

mathslover (mathslover):

so it is given that .. 32 more girls join the group that is x+32 ... r u getting this @Cedrick470 ?

mathslover (mathslover):

increase % = increase / initial amount * 100 %

OpenStudy (anonymous):

oh i hated these they sucked xD

mathslover (mathslover):

u can do it further .. @cedrick470

OpenStudy (anonymous):

Let b and g be the number of boys and girls in the club respectively. Solve the following equations for b and g.\[\left\{b+g=160,\frac{g+32}{160+32}=\frac{g}{160}+\frac{5}{100}\right\}\]\[\{b=48,g=112\} \]